A 1.5 N stone is thrown vertically upward with an initial speed of 15 m/s. Calculate



a. P.E at the maximum height


b. K.E at the instant it will strike the ground


c. P.E after 1 s


d. K.E after 1 s



1
Expert's answer
2022-03-31T11:48:46-0400

(a)

"PE_{max}=mgh_{max}=\frac{mv_0^2}{2}\\
=\frac{1.5/9.8*15^2}{2}=17\:\rm J"

(b)

"KE_f=KE_i=\frac{mv_0^2}{2}\\
=\frac{1.5/9.8*15^2}{2}=17\:\rm J"

(c)

"PE=mgh=mg(v_0t-gt^2/2)\\
=1.5(15*1-9.8*1^2/2)=15\:\rm J"

(d)

"KE=KE_i-PE\\
=17\:\rm J-15\: J=2\: J"